OpenStudy (anonymous):

solve equation 5(4x - 5) = x + 4

OpenStudy (anonymous):

x = 29/19

OpenStudy (mimi_x3):

20x -25 = x + 4 19x = 29 x = 29/19

OpenStudy (anonymous):

19x=29 x=1 10/19 or 29/19